### v1

Abstract
This study investigates the timing and spectral characteristics of X-ray bursts from the neutron star system EXO 0748-676 using the NuSTAR observatory’s FPMA and FPMB instruments in the 3-79 keV range. We identify Type I X-ray bursts driven by thermonuclear explosions on the neutron star’s surface, notably a significant burst at , indicating rapid energy release, followed by a recoil burst at , reflecting stabilization. The correlation between burst timing and the neutron star’s optical period suggests modulation by its rotation and periodic accretion dynamics. Spectral modeling reveals a photon index of and a cutoff energy of , indicating a hot corona around the neutron star. The measured flux of approximately underscores the dynamic nature of accretion-driven systems. Calculated luminosities derived from distance estimates range from to . Comparative analysis with prior observations from the IBIS/ISGRI instrument on the INTEGRAL satellite shows variability in emission characteristics, including softer photon indices and higher cutoff energies in 2003 and 2004. Our examination of smaller energy gaps (3-7 keV, 7-12 keV, etc.) reveals energy-dependent behavior in burst characteristics, enhancing our understanding of nuclear burning phases. Overall, these findings validate models describing Type I X-ray bursts and lay the groundwork for future investigations into similar astrophysical systems and stellar evolution processes in extreme environments.
Keywords NuSTAR accretion accretion discs-stars neutron-pulsars individual EXO 0748-676 X-rays low mass x-ray binaries (LMXBs)
1 Introduction
The idea of low-mass X-ray binaries (LMXBs) Liu et al. (2001) emerged in the late 1960s and early 1970s, soon after X-ray sources were discovered in space. Research on binary systems involving compact objects like neutron stars or black holes began with the identification of bright X-ray sources by satellites such as Uhuru, the first satellite dedicated to X-ray observations, launched in 1970 Pioneer (2023).
The groundbreaking work that identified X-ray binaries as a unique category of astronomical objects is largely attributed to Riccardo Giacconi, Herbert Gursky, and their colleagues. Giacconi, who received the Nobel Prize in Physics in 2002, was instrumental in both the discovery of X-ray sources and the advancement of X-ray astronomy.Additionally, high-mass X-ray binaries (HMXBs) produce strong and variable X-ray emissions due to interactions with the winds from their massive companion stars Charles and Coe (2003), while low-mass X-ray binaries (LMXBs)Liu et al. (2001) usually exhibit more stable X-ray emissions because the material transfer from their smaller companion stars is more consistent. Studying both types of binary systems is important because findings from one can help us understand the other, providing a clearer picture of how these binary star systems evolve over time Kahabka and Van Den Heuvel (1997).
EXO 0748–676 is a well-known neutron star (NS) in a low-mass X-ray binary (LMXB) system that has garnered attention for its frequent and diverse thermonuclear (type I) X-ray bursts. Discovered by the European X-ray Observatory Satellite (EXOSAT), EXO 0748-676 in February 1985 Parmar et al. (1986). EXO 0748-676 is recognized for producing both type I X-ray bursts, caused by thermonuclear explosions on the neutron star’s surface, and continuous X-ray emissions from the accretion process Gottwald et al. (1987). These phenomena provide valuable insights into the behavior of matter under extreme conditions. several studies have focused on the binary system EXO 0748-676, utilizing NuSTAR Harrison et al. (2013) and XMM-Newton Bhattacharya and Bhattacharyya (2024) to better understand its complex behaviour. Recent NuSTAR observations in 2024 captured high-energy phenomena, such as Type I X-ray bursts and consecutive eclipses, shedding light on the accretion processes around the neutron star and the system’s post-quiescence evolution. In parallel, other research combined NuSTAR’s data with XMM-Newton’s high-resolution spectroscopy Bhattacharya and Bhattacharyya (2024) to analyze changes in emission lines before, during, and after bursts. These studies highlighted how circumstellar material and physical processes evolve during outbursts, deepening insights into the neutron star’s environment. Together, these findings enhance our understanding of the system’s thermonuclear bursts and eclipsing features.
EXO 0748-676 is notable for being the only source where gravitationally redshifted absorption lines during X-ray bursts have been observed Cottam et al. (2002); Luo et al. (2013). Identified redshifted spectral lines corresponding to O and Fe transitions during X-ray bursts from EXO 0748-676 ,using the Reflection Grating Spectrometer(RGS) instrument on XMM-newton, indicated a gravitational redshift of z=0.35 at the neutron star’s surface, corresponding to a mass-to-radius ratio of . This finding provides an important empirical constraint on the equation of state (EoS) for dense, cold nuclear matter Cottam et al. (2002).
In 1985, EXO 0748-676 was found to show dips and eclipses every 3.82 hours Parmar et al. (1986); Southwell et al. (1996); Degenaar et al. (2014), with each eclipse lasting 8.3 minutes Psaradaki et al. (2018). These periodic X-ray dips and eclipses were detected by EXOSAT, indicating that material is regularly blocking the X-ray source, likely due to interactions between the companion star and the accretion disk around the neutron star. These features suggest that EXO 0748-676 is viewed at an angle of about Parmar et al. (1986), almost aligned with the plane of the accretion disk. There are several low-mass X-ray binaries (LMXBs)Liu et al. (2001) with high inclinations, but only a few of them exhibit eclipses. Notable examples include EXO 0748-676 Parmar et al. (1986) and MXB 1659-298Sidoli et al. (2001).
Due to the high inclination of EXO 0748-676, the observer’s viewpoint is nearly aligned with the accretion disk, making it ideal for analyzing structures above the disk via spectral lines. High-resolution spectra from Chandra/HETG Jimenez-Garate et al. (2003) and XMM-Newton/RGS have been previously used to study this system. Cottam Cottam et al. (2001)compared spectra during different phases, such as low-emission periods, rapid variations, and bursts, detecting broadened recombination lines from nitrogen, oxygen, and neon. Additionally, T.M.S. Kallman and his team found a gravitationally redshifted absorption feature (z = 0.35) in the X-ray spectrum of EXO 0748-676 Cottam et al. (2002), which is important for determining key neutron star properties, like the ratio of its radius-to-mass Bhattacharyya et al. (2006) and exploring the superdense matter within the neutron star core Bhattacharyya (2010).
This study thoroughly investigates EXO 0748-676, a neutron star located in a Type I X-ray binary system, intending to uncover its distinct features through detailed timing and spectral analyses. Data from NuSTAR observations were meticulously analyzed to gain insights into the unique behavior and properties of the source. The paper is structured as follows: the introduction in section 1 provides an overview of the study and its significance; section 2 explains the data reduction process from NuSTAR; section 3 presents the results, which are further divided into two subsections: timing analysis in subsection 3.1 and spectral analysis in subsection 3.2. The conclusion is summarized in section 4, followed by data availability information in section 5 and a list of references. This structure allows readers to navigate the study easily and understand the various components of the investigation.
2 Observation and data reduction
NuSTAR (Nuclear Spectroscopic Telescope Array), launched by NASA in 2012, Harrison et al. (2013) is the first telescope capable of focusing high-energy X-rays (3-79 keV). It provides unprecedented clarity for studying energetic events such as black holes, neutron stars, and supernova remnants. NuSTAR’s advanced technology uses X-ray mirrors and detectors to capture detailed images of high-energy environments, complementing other observatories like Chandra Degenaar et al. (2009) and XMM-Newton Bhattacharya and Bhattacharyya (2024). Its significant contributions include detecting X-ray jets from black holes, studying supernova remnants, and deepening our understanding of the physics around compact objects like neutron stars. NuSTAR (Nuclear Spectroscopic Telescope Array) recently observed EXO 0748-676 on June 17, 2024, recording an average net count rate of 5.0 counts per second (in the 3-79 keV energy range, with a 2 arcminute extraction radius and both focal plane modules combined).
The lightcurve for the source EXO 0748-676 is shown in Fig. 2, while the detailed information regarding the observation IDs and their respective exposure times is provided in Table 1.
The data processing for EXO 0748-676 was carried out using HEA-SOFT 6.34 includes XSPEC v12.14.1 and CALDB v20240930 for NuSTAR. Light curves, spectra, response matrices, and effective area files were generated with NUSTARDAS software v2.1.4. Unfiltered event files were cleaned using the mission-specific tool NUPIPELINE. A circular source region of 100" was defined around the source center, with a background region selected away from the source (as shown in Fig. 1). Light curves and spectra were extracted from the defined source and background regions using XSELECT and NUPRODUCTS commands. Background subtraction was handled separately for FPMA and FPMB data, and their light curves were combined using the LCMATH tool. The NuSTAR observations of the system were then analyzed for timing and spectral data.
Observation | Observation Start Date | Exposure Time |
ID | Time(UT) | (ks) |
91001324002 | 2024-06-17 01:31:09 | 62.648 |
3 Data analysis and Results
3.1 Timing analysis
In studying the NuSTAR observatory data of the low-mass X-ray binary (LMXB) Liu et al. (2001) EXO 0748-676, after obtaining the time curve lc (light curve) file and the spectral data, PHA (Pulse Height Amplitude) file using the nuproducts command, the background noise was subtracted from the source using the lcmath command. This background subtraction was carried out separately for both the FPMA and FPMB observation data. Once the background was successfully removed from each observation, the two source data sets (FPMA and FPMB) were combined into a single file using lcmath. Afterward, barycentric correction was applied using the barycorr command along with the appropriate auxiliary file containing the necessary information for the correction. This process corrected the timing data to account for the Earth’s motion, adjusting the light curve to the Solar System barycenter for accurate pulsar timing analysis.
After completing the above steps, the light curve was plotted by selecting the relevant time intervals and applying binning to improve the time resolution, using the lcurve command. The light curve was then divided into different energy bands to analyze energy-dependent variations in the count rate. Energy filters corresponding to specific ranges (e.g., 3–7 keV, 7–12 keV, etc.) were applied to split the light curve into segments, allowing for a detailed study of the source’s behavior across different energy ranges and detecting burst events. This was done using the nuproducts command, where specific energy ranges were defined in terms of PI channels, followed by the same steps as above using lcmath for background subtraction and barycorr for barycentric correction for all the energy bins. In Fig. 2, the time curve for the full energy range (3–79 keV) is shown alongside smaller energy ranges (3–7 keV, 7–12 keV, up to 30–79 keV) to examine the nature of the outburst in EXO 0748-676, enabling a detailed analysis of how the outburst evolves across different energy bands.
The analysis of X-ray bursts in the 3-79 keV energy range from the neutron star system reveals significant characteristics of Type I X-ray bursts Gottwald et al. (1987), commonly observed in low-mass X-ray binaries (LMXBs). These bursts are thermonuclear explosions resulting from the unstable burning of accreted material, predominantly hydrogen and helium, on the surface of the neutron star. Such bursts exhibit rapid rise times (0.5-5 seconds) followed by slower decays (10-100 seconds) Strohmayer and Bildsten (2003), releasing immense amounts of energy—on the order of ergs in just a few seconds. This energy output is vastly greater than the Sun’s weekly energy release, driven by runaway nuclear reactions ignited when sufficient material accumulates on the neutron star’s surface Bhattacharyya (2009).
The first notable burst in this dataset occurs at , with a count rate of 28.5, indicating a modest energy release. This is likely a result of localized instabilities in the neutron star’s magnetosphere or fluctuations in its accretion flow, producing only a minor deviation in the X-ray count rate. Such smaller bursts are typical of low-level accretion-driven instabilities but do not represent major thermonuclear events.
A dramatic shift in behavior is observed at , where a major burst is detected with a count rate of 1,615.88. This burst is indicative of a powerful Type I X-ray burst, caused by thermonuclear fusion on the neutron star’s surface Galloway et al. (2008a). Type I bursts occur when accreted material from a companion star accumulates on the neutron star’s surface, reaching critical density and temperature. At this point, a runaway nuclear reaction is triggered, releasing vast amounts of energy in a short span. This particular burst likely involved helium ignition, where the accreted matter had built up over hours or days, reaching a column density of approximately and igniting through the unstable triple-alpha process Bhattacharyya (2009). Such bursts are frequently observed in LMXBs and are marked by sudden, intense X-ray emissions as the surface material ignites in a thermonuclear explosion Liu et al. (2001).
This analysis of the X-ray burst data in the 3-79 energy range, combined with the neutron star’s optical period of 13,750 seconds Parmar et al. (1986); Knight et al. (2022); Southwell et al. (1996), reveals key aspects of the system’s behavior. It supports the interpretation that the burst represents a significant thermonuclear event, while the smaller bursts, such as the one at , likely result from less energetic fluctuations related to the star’s accretion system or magnetospheric activity. These findings align with the well-documented behaviors of Type I X-ray bursts, where accretion, nuclear burning, and rotational dynamics all interplay to generate recurrent, explosive energy releases.
Following the main burst at , another burst is detected at with a count rate of 32.14. This event likely represents a "recoil" burst, where the system releases residual energy after the primary thermonuclear explosion. It is common in neutron star systems for smaller, secondary bursts to follow the main burst as the system stabilizes. These recoil bursts are driven by remaining instabilities in the accreted material or disturbances in the neutron star’s magnetic field, allowing the system to return to equilibrium.
Subsequent smaller bursts occur at and , with count rates of 32.93 and 27.55, respectively. These bursts, although significant, are much less energetic than the primary event. They likely result from local fluctuations in the neutron star’s accretion flow or minor magnetic reconnections. These types of events are characteristic of accreting neutron stars, where variations in material flow and magnetic activity can produce periodic energy releases in the form of X-ray emissions.
Overall, these secondary bursts, including the recoil event, reflect the system’s post-burst behavior, where residual energy dissipation and localized adjustments in the accretion process or magnetosphere lead to smaller bursts of energy. While not as dramatic as the main thermonuclear explosion, these events are still important in understanding the complex dynamics of neutron star systems.
The intervals between these bursts provide further insights into the system’s dynamics. The gap between the first major burst at and the recoil burst at is approximately 984 seconds, reflecting rapid energy release and adjustment. More intriguingly, the gap between the burst at and the next significant burst at is roughly 13,778 seconds, closely matching the neutron star’s optical period of 13,750 seconds. This suggests a potential link to the star’s rotational cycle, where some bursts could be influenced by periodic accretion events associated with its binary interaction, pointing to the role of rotational dynamics and the magnetic field structure in modulating the burst mechanism.
Supporting the thermonuclear origin of these bursts, studies by Strohmayer et. al. Strohmayer (2006); Strohmayer and Bildsten (2003) demonstrate that the burst emission area matches the expected neutron star surface area, confirming that the bursts stem from nuclear burning rather than accretion disk processes. Earlier theoretical predictions by Hansen and Van Horn (1975) Hansen and Van Horn (1975) suggested that neutron stars with thin hydrogen and helium layers would undergo such unstable nuclear burning, as stable burning would result in energy dissipation through gravitational processes.
These bursts follow distinct burning regimes based on the accretion rate. At moderate accretion rates, hydrogen burns steadily via the hot CNO cycle Strohmayer and Bildsten (2003); Fowler and Hoyle (1965); Bildsten (2000), allowing helium to accumulate and ignite in a sudden burst, referred to as a helium burst. In cases where the burst surpasses the Eddington luminosity, a photospheric radius expansion (PRE) burst may occur, where radiation pressure lifts the neutron star’s outer layers. In addition to the well-documented Type I X-ray bursts, another class of thermonuclear bursts known as superbursts has been observed. These bursts are characterized by significantly longer recurrence times, on the order of years, and differ from Type I bursts in both decay time and energy output. While Type I bursts last for seconds to minutes, superbursts exhibit much slower decay times, ranging from 1 to 3 hours, and release vastly more energy, around ergs. The first superburst was discovered by Cornelisse et al. Cornelisse et al. (2000), marking a major distinction from the shorter and less energetic Type I bursts typically seen in neutron star systems.
The energy range over which X-ray bursts are observed plays a significant role in understanding the physical mechanisms driving these thermonuclear explosions on neutron stars. In the broad 3-79 keV energy range, the burst at stands out as the most energetic event, with a count rate of 1615.88 counts per second. This range captures the full evolution of the burst, from the initial thermonuclear ignition to the cooling phase, providing a complete picture of the burst dynamics. The broad energy spectrum shows not only the intense peak but also the subsequent smaller bursts, such as the recoil burst at , with a count rate of 32.14. These secondary bursts likely result from residual energy dissipation or magnetospheric adjustments. However, when we break down the burst behavior into narrower energy ranges, important patterns emerge. In the soft X-ray band (3-7 keV), the primary burst shows a much lower count rate of 647.47 counts per second, indicating that much of the burst’s energy is emitted in higher energy bands. Soft X-rays primarily capture the cooling phase of the burst, which decays more slowly, highlighting the prolonged emission from the neutron star’s surface. This behavior aligns with our understanding of Type I X-ray bursts, where initial high-energy photons rapidly decline, leaving softer X-rays to dominate the emission for longer periods Galloway et al. (2008b).
In contrast, mid-energy ranges (7-12 keV and 12-18 keV) show bursts with intermediate intensity, emphasizing the importance of these bands in capturing the peak of the thermonuclear explosion. For instance, in the 7-12 keV range, the primary burst reaches 752.16 counts per second, while in the 12-18 keV range, it drops to 189.09 counts per second. This suggests that the bulk of the burst’s energy is emitted in these mid-energy bands, where nuclear burning on the neutron star’s surface releases a significant amount of X-rays before cooling. The physics here is driven by helium burning, which ignites under conditions of high pressure and temperature on the star’s surface, leading to rapid energy release followed by radiative cooling Strohmayer and Bildsten (2003); Fowler and Hoyle (1965); Lewin et al. (1997, 1993).
Higher energy bands (18-30 keV and 30-79 keV) capture the hardest X-rays, but the intensity is much lower compared to the full energy range. For instance, in the 18-30 keV band, the primary burst has a count rate of 36.57 counts per second, while in the 30-79 keV band, it further decreases to 6 counts per second. These hard X-rays are likely emitted during the early phases of the burst, reflecting the hottest parts of the thermonuclear explosion. The rapid decline in higher energy bands supports the idea that hard X-rays are primarily linked to the initial stages of nuclear burning, with less emission as the system cools. Such behavior is consistent with the high-energy photons emitted during the peak of a photospheric radius expansion (PRE) burst, where radiation pressure lifts the outer layers of the neutron star before settling into a cooling phase Strohmayer (2006); Strohmayer and Bildsten (2003).
3.2 Phase-averaged Spectral analysis
The X-ray spectral analysis of the source observed by NuSTAR’s FPMA and FPMB instruments across the 3-79 keV energy range provides critical insights into the emission processes of this system. The analysis was performed using the xspec tool (v12.14.1) Arnaud (1996) to model the observed spectrum, which was initially grouped to ensure a minimum of 20 counts per spectral bin using the ftool GRPPHA. This binning process improves the statistical significance of the spectral data by minimizing the impact of Poisson noise, particularly in regions with low photon counts. To account for the relative calibration differences between FPMA and FPMB, a constant model was applied, fixing FPMA to unity and allowing the FPMB normalization to vary.
Spectral parameters | NuSTAR(2024) | IBIS | |
ISGRI(2003) | ISGRI(2004) | ||
6.350.15 | - | - | |
Photon Index | 1.240.014 | 1.30.4 | 1.60.4 |
Cutoff Energy (keV) | 36.201.04 | 4415 | 5015 |
1.77E-024.39E-04 | - | - | |
1864.80 | - | - | |
Degrees of freedom | 1899 | - | - |
0.98 | - | - | |
Flux | 381.170.016 | 320150 | 421150 |
The observed spectrum was fitted using a model combination of CONSTANT PHABS CUTOFFPL, which includes three key components. The PHABS model accounts for photoelectric absorption due to neutral interstellar matter, characterized by the hydrogen column density (). The CUTOFFPL model is a power-law with an exponential cutoff, typically used to describe the X-ray emission in accreting systems, especially those containing compact objects like neutron stars. The spectral fit resulted in reduced chi-squared values () of 0.98 for FPMA and 1.007 for FPMB, indicating that the model accurately describes the data, as illustrated in Fig. 3.
The fit revealed several key parameters. The hydrogen column density was found to be , suggesting significant absorption along the line of sight, likely due to dense interstellar matter or local absorbing material. The photon index () indicates that the X-ray spectrum is relatively hard, pointing to non-thermal processes or Comptonized thermal photons in a hot corona around the neutron star. The cutoff energy () reflects the energy at which the X-ray emission begins to decline sharply, possibly due to thermal Comptonization processes in the corona. The flux in the 3-79 keV energy range was measured at , providing a useful reference for future observations.
These results were compared with previous observations by the IBIS/ISGRI instrument on the INTEGRAL satellite, which observed the source in the 20-100 keV range in 2003 and 2004. The earlier IBIS/ISGRI data showed slightly softer photon indices, with values of (2003) and (2004), suggesting the emission was somewhat softer during those observations. Additionally, the cutoff energies were higher, 44 keV in 2003 and 50 keV in 2004, compared to NuSTAR’s 36.2 keV, indicating a potentially hotter Comptonizing region or corona in those periods. The flux values also varied between in 2003 and in 2004, reflecting typical fluctuations in accretion-driven systems. The ISGRI detector, designed for hard X-ray to soft gamma-ray observations, is well-suited for studying sources like EXO 0748-676 during both outburst and quiescent phases. Their spectral fitting was performed using XSPEC v11.2, and a 10% systematic error was added to account for uncertainties in their spectral extraction and response matrix. The fitting was done using a cutoff power-law model, where the photon index () and cutoff energy () parameters matched the general trends seen in NuSTAR’s observations.
Based on the gravitational redshift, Özel suggested that the distance of EXO 0748-676 is 9.2 1.0 kpc Özel (2006). From a strong X-ray burst, Wolff et al. Wolff et al. (2005) derived a distance of 7.7 kpc for a helium burst. Galloway et al. analyzed several type I X-ray bursts from EXO 0748-676 and estimated a distance of 7.4 kpc Galloway et al. (2008b), different from the value of 9.2 kpc reported by Özel Özel (2006). Taking into account the touchdown flux and the high binary inclination of EXO 0748-676, Galloway, Özel, and Psaltis recently gave a distance of 7.1 1.2 kpc Galloway et al. (2008a); Zhang et al. (2011).
Using a calculated flux of , the corresponding luminosities for these distances were derived. For a distance of 9.2 kpc, the luminosity is . At a distance of 7.7 kpc, the luminosity is . For a distance of 7.4 kpc, the luminosity is , with a negligible error due to the exact distance value. Finally, for a distance of 7.1 kpc, the calculated luminosity is . These results demonstrate the relationship between distance and luminosity, with greater distances resulting in higher luminosity values.
The luminosities derived from different distance estimates of EXO 0748-676 in our study can be compared to those reported in Trigo et al. (2006); Boirin et al. (2007). Using the data from revolution 719, Díaz Trigo et al. Trigo et al. (2006) report an absorbed flux in the range of 0.6–10 keV measured at . This absorbed flux, when corrected for no attenuation, corresponds to an unabsorbed flux of . Boirin et al. Boirin et al. (2007) further calculated an unabsorbed bolometric flux of across the 0.1-100 keV energy band. At a distance of 5 kpc, this flux corresponds to a luminosity of , while at 10 kpc, it yields a significantly higher luminosity of .
In contrast, our analysis using the calculated flux of erg/cm²/s provides a range of luminosities based on various distance estimates. For a distance of 9.2 kpc, the luminosity is erg/sec, and at 7.7 kpc, it is erg/sec. As the distance decreases to 7.4 kpc, we find a luminosity of erg/sec, and at 7.1 kpc, the luminosity is erg/sec.
When comparing these results with Díaz Trigo’s estimates of at 5 kpc and at 10 kpc in the 0.6-10 keV range, we observe that our findings are in line with these estimates, particularly when considering the range of distances. Specifically, our luminosity results suggest that the pulsar is relatively bright at the higher distance of 9.2 kpc, which aligns with the higher end of the luminosities reported by Díaz Trigo. Additionally, the decreasing luminosities associated with shorter distances emphasize the inverse relationship between distance and brightness.
These results also underscore the critical role of accurate distance measurements in astrophysical studies, as variations in distance can significantly affect luminosity calculations. The discrepancies between our luminosity estimates and those reported by Boirin et al. at different distances may stem from differences in energy ranges analyzed or variations in the flux values used in their calculations. Furthermore, our findings highlight the importance of using multiple distance estimates to obtain a comprehensive understanding of the source’s intrinsic properties, as the luminosity values can provide insights into the physical processes occurring in EXO 0748-676, particularly during burst events. Overall, the luminosity derived from our analysis demonstrates that EXO 0748-676 remains a key target for understanding the behavior of low-mass X-ray binaries, contributing to the ongoing discourse regarding the physical mechanisms at play during X-ray outbursts.
4 Conclusion
In conclusion, the timing analysis of X-ray bursts from the neutron star system provides valuable insights into the interplay of thermonuclear processes, accretion dynamics, and magnetic field interactions. The primary burst at signifies a substantial thermonuclear explosion on the neutron star’s surface, marked by a sharp increase in count rate, indicating a rapid energy release. Following this major burst, smaller events, such as the recoil burst at , illustrate the system’s stabilization after the initial explosion. The alignment of burst timings with the neutron star’s optical period suggests a possible modulation by the star’s rotation and periodic accretion, revealing how these factors influence the timing and intensity of bursts. Additionally, the analysis across different energy bands highlights the relationship between energy levels and burst characteristics, with soft X-rays linked to cooling phases, mid-energy emissions representing peak explosions, and hard X-rays associated with thermonuclear ignition. Overall, these findings reinforce our understanding of the mechanisms behind Type I X-ray bursts and pave the way for future research into the physics governing accretion and burst phenomena in neutron star systems, enhancing our knowledge of stellar evolution and the extreme conditions in the universe.
The spectral analysis of EXO 0748-676 conducted using the NuSTAR observatory’s FPMA and FPMB instruments in the 3-79 keV energy range provides significant insights into the source’s emission characteristics, revealing a photon index of and a cutoff energy of , which suggests the presence of a hot corona around the neutron star and indicates the nature of the emission processes at play. The measured flux of approximately highlights the variable nature of accretion-driven systems. Comparative analysis with earlier observations from the IBIS/ISGRI instrument on the INTEGRAL satellite further enhances our understanding of the source’s behavior over time, as the slightly softer photon indices and higher cutoff energies observed in 2003 and 2004 suggest that the emission may have been influenced by changes in coronal temperature and accretion dynamics during those periods. The variability in flux measurements across the different observational epochs underscores the dynamic nature of the source, reflecting typical fluctuations seen in accreting neutron stars. Overall, this analysis illustrates the effectiveness of high-energy spectral fitting in probing the physical processes governing the emission from EXO 0748-676, contributing to the broader understanding of neutron star physics and the complexities of accretion in binary systems, while the consistency of spectral parameters with previous studies also validates the reliability of the methods employed, establishing a foundation for future investigations into similar astrophysical objects.
Additionally, our analysis of the luminosity of EXO 0748-676, derived from various distance estimates and flux measurements, provides critical insights into the characteristics of this low-mass X-ray binary. The calculated luminosities vary based on distance, highlighting the importance of accurate distance measurements in astrophysical studies. Notably, our results align well with previous estimates from Díaz Trigo and Boirin et al., suggesting that EXO 0748-676 exhibits significant intrinsic brightness, particularly at the higher distance of 9.2 kpc. These findings emphasize the need for continued investigation into the physical processes governing X-ray bursts in this system, as understanding the luminosity and behavior of such sources is essential for elucidating the mechanisms driving X-ray emission in low-mass X-ray binaries. Ultimately, this study contributes to a more comprehensive understanding of EXO 0748-676 and its role in the broader context of X-ray astronomy.
5 Data Availability
The data utilized in this study is available to the public through the HEASARC data archive, allowing for research purposes.